Math939 Fall 2019

References

[1] Alligood, Kathleen T., Tim D. Sauer, and James A. Yorke. ”Chaos: An Introduction to Dynamical Systems. 1996.” (1997).

[2] Hale, Jack K., and Hüseyin Koçak. Dynamics and bifurcations. Vol. 3. Springer Science & Business Media, 2012.

[3] Matsumoto, T., Komuro, M., Kokubu, H., Tokunaga, R., Bifurcations – Sights, Sounds, and Mathematics, Springer-Verlag, 1992.

[4] Robinson, Clark. Dynamical systems: stability, symbolic dynamics, and chaos. CRC press, 1998.

[5] Shub, Michael. Global stability of dynamical systems. Springer Science & Business Media, 2013.